OpenStudy (anonymous):

4x+3y=55 and 5x-3y=35 elimination method

OpenStudy (anonymous):

add the two equations 4x+3y= 55 +5x -3y = 35 ---------------- 9x=90 x = 10

OpenStudy (anonymous):

i did that and i got x rigth but finding why isnt working

OpenStudy (anonymous):

finding y*

OpenStudy (anonymous):

u got x as 10..now put in x as 10 in any one eqn and find y

OpenStudy (anonymous):

can you show how please

OpenStudy (anonymous):

please

OpenStudy (anonymous):

see x=10 4(10) + 3y = 55 40+3y=55 3y=15 y=5

OpenStudy (anonymous):

where did u get 15 from

OpenStudy (anonymous):

40 + 3y = 55 40 + 3y - 40 = 55-40 40-40 + 3y = 55-40 3y=15 y=5